# Training robust and generalizable quantum models

### v3

@@ C.1 Numerical setup
All numerical simulations within this work were performed using the Python QML library PennyLane [61]. As device, we used the noiseless simulator "lightning.qubit" together with the adjoint differentiation method, to enable fast and memory efficient gradient computations. In order to solve the optimization problem in (9), we apply the ADAM optimizer using a learning rate of $\eta=0.1$ and the suggested values for all other hyperparameters [62]. Furthermore, we run $200$ epochs throughout and train $12$ models based on different initial parameters for varying regularization parameters $\lambda\geq 0$. Adding the regularization does not introduce significant computational overhead as the evaluation of the cost only involves a weighted sum of the terms $\|w_{j}\|^{2}$. As final model, we take the set of parameters for the model with minimal cost over all runs and epochs. Furthermore, the training as well as the robustness and generalization analysis were parallelized using Dask [63].
@@
In our numerical case study we set $\gamma_{j}=0$ for all $j$ since this 1) still allows to reach arbitrary points on the Bloch sphere and 2) enables an easier comparison to fixed-encoding quantum models. In order to introduce entanglement in a hardware-efficient way, we use a ring of CNOTs. The considered circuit is shown in Figure 7 and involves three layers of rotations and entanglement, which we observed to be a good trade-off between expressivity and generalization. More precisely, in our simulations, fewer layers were not sufficient to accurately solve the classification task, whereas more layers led to higher degrees of overfitting. For the fixed-encoding quantum models as in (14), we use a similar $3$-layer ansatz, encoding the two entries of the 2D data points into the first and second angle $\alpha_{j}$, $\beta_{j}$ of the rotation gates, followed by a parametrized $\mathrm{SU(2)}$ rotation with three free parameters. We also perform a classical data pre processing, scaling the input domain from $[-1,+1]$ to $[-\pi,+\pi]$, such that the full possible range of the rotation angles can be utilized.

### v4

@@ C.1 Numerical setup
All numerical simulations within this work were performed using the Python QML library PennyLane [62]. As device, we used the noiseless simulator ”lightning.qubit” together with the adjoint differentiation method, to enable fast and memory efficient gradient computations. In order to solve the optimization problem in (9), we apply the ADAM optimizer using a learning rate of $\eta=0.1$ and the suggested values for all other hyperparameters [63]. Furthermore, we run $200$ epochs throughout and train $12$ models based on different initial parameters for varying regularization parameters $\lambda\geq 0$. Adding the regularization does not introduce significant computational overhead as the evaluation of the cost only involves a weighted sum of the terms $\|w_{j}\|^{2}$. As final model, we take the set of parameters for the model with minimal cost over all runs and epochs. Furthermore, the training as well as the robustness and generalization analysis were parallelized using Dask [64].
@@
In our numerical case study we set $\gamma_{j}=0$ for all $j$ since this 1) still allows to reach arbitrary points on the Bloch sphere and 2) enables an easier comparison to fixed-encoding quantum models. In order to introduce entanglement in a hardware-efficient way, we use a ring of CNOTs. The considered circuit is shown in Figure 7 and involves three layers of rotations and entanglement, which we observed to be a good trade-off between expressivity and generalization. More precisely, in our simulations, fewer layers were not sufficient to accurately solve the classification task, whereas more layers led to higher degrees of overfitting. For the fixed-encoding quantum models as in (14), we use a similar $3$-layer ansatz, encoding the two entries of the 2D data points into the first and second angle $\alpha_{j}$, $\beta_{j}$ of the rotation gates, followed by a parametrized $\mathrm{SU(2)}$ rotation with three free parameters that are optimized, compare $\phi_{j}$ in (11). We also perform a classical data pre processing, scaling the input domain from [1,+1][-1,+1] to $\pi,+\pi]$, such that the full possible range of the rotation angles can be utilized.